# FANformer: Improving Large Language Models Through Effective Periodicity Modeling

### v1

@@ 4.1 Implementation Details
StableLM-base-alpha-v2-3B 3B 1T 32.4 67.3 64.6 68.6 26.4 76 89.5 62.1 60.9
TinyLlama-v1.1-1T 1.1B 1T 33.1 49.5 58.4 52.5 37.8 70.4 86.4 55.2 55.4
@@
4.2 Scalability of FANformer
@@
We follow OLMo’s configuration and vary the FFN’s intermediate dimension $d_{f}$ to keep the number of parameters consistent for all models in this experiment. For scaling up mode parameters, we adopt Dolma v1_6-sample as training data and train LLMs from 268M to 7B. We compare FANformer with the standard Transformer and a variant of FANformer, termed Transformer+ATM, which uses MLP layer instead of FAN layer in FANformer. For scaling up training tokens, we train 1B LLMs on the first 200 billion of our sampled 1T tokens.
Figure 3: Language modeling loss of scaling up model parameter and training tokens. Left: we train LLMs from 268M to 7B parameters. Right: we evaluate 1.0B LLMs every 20B tokens up to 200B tokens.
Results.
As shown in Figure 3, the scaling law (Kaplan et al., 2020) empirically aligns well with the results obtained from our FANformer, underscoring its superior scalability properties. Figure 3 (left) reveals that the implementation of FAN consistently surpasses the performance of the standard Transformer across a range of model sizes. This finding highlights FAN’s enhanced scalability in terms of parameter efficiency, as it achieves comparable performance with only 69.2% of the parameters required by the standard Transformer. Notably, the scaling curve of Transformer+ATM closely overlaps with that of the standard Transformer, indicating that merely revising attention mechanisms using MLP Layer is insufficient. This observation further emphasizes that FANformer’s performance gains are not attributable to network depth increase, but rather to its special architectural design. Figure 3 (right) demonstrates that FANformer achieves performance parity with the standard Transformer while utilizing significantly fewer training tokens. Specifically, FANformer requires only 159.6B training tokens to match the performance of the standard Transformer trained on 200B tokens, representing a 20.3% reduction in training resource requirements. These findings suggest that FANformer exhibits superior utilization efficiency in terms of both model parameters and training tokens compared to the standard Transformer architecture.
4.3 Performance of FANformer-1B
We pretrain FANformer-1B on 1 trillion tokens and report zero-shot performance on a set of commonsense downstream tasks, following previous work (Brown et al., 2020; Touvron et al., 2023a; Groeneveld et al., 2024, inter alia).
Setup.
The downstream evaluation suite consists of 8 core tasks, including ARC-C (Clark et al., 2018), ARC-E (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA Mihaylov et al. (2018), PIQA (Bisk et al., 2020), SCIQ (Welbl et al., 2017), and WinoGrande (Sakaguchi et al., 2020). We compare pretrained FANformer-1B to seven open-source LLMs with comparable model parameters or training tokens, including Qwen2.5-1.5B (Team, 2024), R1-Distill-Qwen1.5B DeepSeek-AI et al. (2025), Llama-3.2-1B Dubey et al. (2024), TinyLlama-v1.1-1B Zhang et al. (2024), OLMo-1B (Groeneveld et al., 2024), OpenLLaMA-v2-3B Geng et al. (2023), and StableLM-base-alpha-v2-3B Tow (2023).
@@ 4.3 Performance of FANformer-1B
FANformer + Activation 1.0×1.0$\times$1.0 × 2.893 3.34 30.64 3.07 24.50 53.61
@@
FANformer + Activation 1.04×1.04$\times$1.04 × 2.887 3.34 30.57 3.07 24.39 53.13
@@ Results.
4.4 Further Analysis
4.4.1 Ablation Study and Variant Analysis
Setup.
We compare FANformer to other variant architectures, including the above-mentioned Transformer+ATM, Transformer+ATL: use two linear transforms to compute QKV projection, FANformer +Activation: employ Eq. (1) with GELU Hendrycks et al. (2016) activation function instead of Eq. (2), Transformer (FFN $\leftarrow$ FAN): replace FFN with FAN Dong et al. (2024b), and standard Transformer as their ablations.
Results.
From Table 4.3, we have the following findings: 1) FANformer consistently outperforms other variant architectures in both scenarios of the same parameter and same dimension on all evaluation metrics. 2) The performance of Transformer+ATM and Transformer+ATL is notably inferior to that of FANformer, indicating that the core improvement stems from the ATF module we designed. 3) Although Transformer (FFN $\leftarrow$ FAN) yields some improvement, this enhancement is inferior to the gains achieved by FANformer, suggesting that integrating periodicity modeling within attention is more advantageous than FFN on language modeling. 4) Incorporating activation functions such as GELU into the attention mechanism tends to degrade model performance. Specifically, FANformer+Activation and Transformer+ATM exhibit weaker performance compared to FANformer and Transformer+ATL, likely because these activation functions suppress certain features, thereby hindering subsequent attention operations.
Figure 4: Training loss of FANformer and Transformer on early training steps. The complete training loss is provided in Figure 8 of Appendix.
@@ 4.4.2 Training Dynamics
We perform a comparative analysis of the loss trends during the training process between our FANformer and Transformer, as illustrated in Figure 4. The experimental results indicate that the loss of FANformer decreases more slowly in the early stages compared to Transformer, which we hypothesize may be due to the initial lack of established periodic modeling. As the training progresses and periodic modeling gradually improves, FANformer demonstrates a faster convergence rate, with its loss decreasing more rapidly than that of Transformer. This result suggests that as the model progressively learns from the data, the learning efficiency of our FANformer notably surpasses the standard Transformer.
@@
Figure 5: Performance of FANformer and Transformer on modular addition and linear regression tasks, where the darkened regions indicate areas where the model performance approaches zero, signifying the emergence of the "hole" as described in the work Hu et al. (2024).
4.4.3 Effect of hyperparameter $p$
We systematically investigate the impact of hyperparameter p, which controls the proportion of periodicity modeling in FANformer, on model performance across its value range. The experimental results from the 1B-scale FANformer (as shown in Figure 6) demonstrate that our model exhibits strong robustness in terms of training loss and downstream task accuracy, with relatively small performance fluctuations. Furthermore, regardless of the variation in p values, FANformer consistently outperforms the standard Transformer (horizontal baseline). Analysis of experimental results from models of different scales (300M, 1B, 3B) (as illustrated in Figure 9 of Appendix) reveals a clear trend: larger models tend to exhibit higher optimal p values. This observation suggests that more powerful FANformers are better equipped to extract more intricate latent periodicity features.
@@ 4.4.4 Case-based and Rule-based Reasoning
Following the work Hu et al. (2024), we evaluate the case-based and rule-based reasoning of Transformer and our FANformer on two tasks, including: (1) Modular addition: $c=(a+b)\mod 113$ with $a,b\in[0,112]$; (2) Linear regression: $c=a+2b+3$ with $a,b\in[0,99]$. We finetune pretained LLMs, i.e., OLMo-1B and FANformer-1B, on each task dataset for 500 epochs and their performance is measured via the Leave-Square-Out method, sampling 10 generations at temperature 0.5 per test point.
@@ Results.
As illustrated in Figure 5, both Tranformer and our FANformer achieve near-perfect accuracy on the training set of modular addition and linear regression tasks, approaching approximately 100%. However, a critical divergence emerges in their performance on the test sets. Specifically, as shown in Figure 5, Transformer exhibits a pronounced failure to generalize, resulting in a "black hole" in the center of the figure, indicating that its accuracy on the test dataset drops to nearly zero. This observation is consistent with the findings reported in the work  (Hu et al., 2024). In contrast, FANformer demonstrates a marked improvement in addressing the "hole" issue. In the linear regression and modular addition tasks, there is no obvious hole observed, further corroborating the hypothesis that, relative to the Transformer-based model, FANformer possesses a stronger tendency to learn underlying rules, thereby achieving superior generalization performance.
@@ 5 Related Work
In this section, we outline the three most relevant directions and associated papers of this work.
5.1 Large Language Models
@@
The rapid advancement of language models has revolutionized natural language processing and artificial intelligence research Radford (2018); Dubey et al. (2024); DeepSeek-AI et al. (2025). The emergence of GPT-3 Brown et al. (2020), with 175B parameters, showcased remarkable few-shot prompting abilities, suggesting that scaling laws Kaplan et al. (2020) could unlock emergent capabilities. Recent notable LLMs like PaLM Chowdhery et al. (2023), LLaMA Touvron et al. (2023a), GPT-4 Chowdhery et al. (2023), and DeepSeek Bi et al. (2024) further pushed the boundaries of model size and performance. Additionally, the open-source release of OLMo Groeneveld et al. (2024) has provided valuable resources for the community, enabling more accessible training of LLMs.
5.2 Advances in Transformer Architecture
Recent advancements in Transformer architecture primarily address two fundamental limitations: computational inefficiency in long-context processing and insufficient expressiveness of attention mechanisms. To tackle long-context processing limitations, Sparse Transformer (Child et al., 2019) and Longformer (Beltagy et al., 2020) introduce sparsity and local Attention. Query mechanism innovations like MQA (Shazeer, 2019) and GQA Ainslie et al. (2023) optimize token interaction efficiency. For inference acceleration, MLA DeepSeek-AI et al. (2024a) pioneers a low-rank latent space compression technique for Key-Value pairs, dramatically reducing cache memory requirements. Hardware-level optimizations emerge through FlashAttention Dao et al. (2022), which strategically minimizes GPU memory access overhead during attention computation. To improve the expressiveness of networks, Probabilistic Attention Keys Nguyen et al. (2022) replace deterministic key-query interactions with Gaussian kernels, enabling better capture of semantic relationships. Selective Attention Leviathan et al. (2024) introduces dynamic token masking to suppress irrelevant features to refine attention mechanism. Differential Transformer Ye et al. (2024) addresses attention noise in long contexts by computing dual softmax maps and performing subtraction.
Different from previous work, we improve language modeling by addressing the challenge of modeling periodicity in Transformers, which can seamlessly incorporate the aforementioned works for revising the attention mechanism, as demonstrated in the derivation provided in Appendix F.
5.3 Fourier-based Neural Networks
@@
Previous research on Fourier-based Neural Networks was aimed at solving some domain-specific applications Zuo et al. (2005); Tan (2006); Chen et al. (2022); Li et al. (2021). Some studies specifically explored the use of sinusoidal activations (e.g., cosine Silvescu (1999) Ngom et al. (2021) or sine Parascandolo et al. (2016); Sitzmann et al. (2020)) to approximate periodic patterns Liu (2013). However, these approaches lacked generalizability beyond narrow domains due to rigid frequency parameterization and limited scalability Uteuliyeva et al. (2020); Liu et al. (2020). FAN (Dong et al., 2024b) introduces Fourier Principle into neural networks to effectively solve periodicity modeling challenges while maintaining broad applicability similar to MLP. FNet Lee-Thorp et al. (2022) replaces self-attention with Fourier Transform to achieve linear complexity, but it sacrifices the performance of language modeling. In contrast, we employ effective periodicity modeling to improve LLMs.
@@ 6 Conclusion
We propose FANformer, a novel LLM architecture that enhances learning efficiency by integrating Fourier Analysis Network into self-attention mechanism for effective periodicity modeling. Experiments demonstrate that FANformer outperforms Transformer when scaling model parameters and training tokens, achieving better performance with 31% fewer parameters and 20% fewer tokens. Pretrained FANformer-1B surpasses open-source LLMs of comparable size or training scale on various downstream tasks, highlighting the benefits of periodicity-aware architecture design. The discovery of FANformer’s enhanced scalability, learning efficiency, and rule-based learning advantages suggests potential pathways for developing more efficient and high-performance language models.
7 Limitations
Second, our work is orthogonal to the existing approaches for revising the attention mechanism, i.e., our work can seamlessly incorporate them, as verified in the derivation provided in Appendix F. There are numerous variants of attention mechanisms, as discussed in the related work (Section 5.2), such as Flash Attention Dao et al. (2022), MQA Shazeer (2019), and MLA DeepSeek-AI et al. (2024a). In this work, we only incorporate Flash Attention for necessary acceleration, while leaving the exploration of other approaches for future work.
@@ Appendix A Training Loss Curves of OLMO and FANformer
Appendix B Extended Results about Hyperparameter $p$ for Section 4.4.3
We investigate the impact of hyperparameter p on training loss across FANformer models of varying scales, with the results illustrated in Figure 9.
@@
Appendix C Instruction Following with SFT
@@ C.2 Evaluation Setup
Appendix D Detailed Results of Ablation Study for Section 4.4.1
In ablation study, we report the average results across various tasks on V2 Validation Sets, V3 Validation Sets, and Downstream tasks, with the specific tasks detailed in Section G.3. The complete results are detailed in Table 4 and Table 5.

### v2

@@ 4.3 Performance of FANformer-1B
StableLM-v2-3B 3B 1T 32.4 67.3 64.6 68.6 26.4 76 89.5 62.1 60.9
TinyLlama-v1.1 (1T) 1.1B 1T 33.1 49.5 58.4 52.5 37.8 70.4 86.4 55.2 55.4
@@
Results.
4.4 Further Analysis
4.4.1 Ablation Study and Variant Analysis
@@
We compare FANformer to other variant architectures, including 1) the above-mentioned Transformer+ATM, 2) Transformer+ATL, which use two linear transforms to compute QKV projection, 3) FANformer (original FAN) that employs Eq. (10) (original FAN layer) instead of Eq. (1) in FANformer, 4) Transformer (FFN $\leftarrow$ FAN) where the FFN is replaced with FAN (Dong et al., 2024a), and 5) standard Transformer as their ablations.
@@
FANformer (original FAN) 1.0×1.0$\times$1.0 × 2.893 3.34 30.64 3.07 24.50 53.61
@@
FANformer (original FAN) 1.04×1.04$\times$1.04 × 2.887 3.34 30.57 3.07 24.39 53.13
@@ Results.
From Table 4.4.1, we have the following findings: 1) FANformer consistently outperforms other variant architectures in both scenarios of the same parameter and same dimension on all evaluation metrics. 2) The performance of Transformer+ATM and Transformer+ATL is notably inferior to that of FANformer, indicating that the core improvement stems from the ATF module we designed. 3) Although Transformer (FFN $\leftarrow$ FAN) yields some improvement, this enhancement is inferior to the gains achieved by FANformer, suggesting that integrating periodicity modeling within attention is more advantageous than FFN on language modeling. 4) Incorporating activation functions such as GELU into the attention mechanism tends to degrade model performance. Specifically, FANformer (original FAN) and Transformer+ATM exhibit weaker performance compared to FANformer and Transformer+ATL, likely because these activation functions suppress certain features, thereby hindering subsequent attention operations.
Figure 4: Training loss of FANformer and Transformer on early training steps, with the complete training loss is provided in Figure 7 of Appendix.
@@ 4.4.2 Training Dynamics
We perform a comparative analysis of the loss trends during the training process between our FANformer and Transformer, as illustrated in Figure 4. The experimental results indicate that the loss of FANformer decreases more slowly in the early stages compared to Transformer, which we hypothesize may be due to the initial lack of established periodic modeling. As the training progresses and periodic modeling gradually improves, FANformer demonstrates a faster convergence rate, with its loss decreasing more rapidly than that of Transformer. Intuitively, once the core of semantic knowledge is established, the model’s inherent periodic modeling lets new concepts attach to existing representations instead of being learned from scratch, accelerating subsequent learning. This result suggests that as the model progressively learns from the data, the learning efficiency of our FANformer notably surpasses the standard Transformer.
@@
Figure 5: Performance of FANformer and Transformer on modular addition and linear regression tasks, where the darkened regions indicate areas where the model performance approaches zero, signifying the emergence of the "hole" as described in the work (Hu et al., 2024).
@@ 4.4.3 Case-based and Rule-based Reasoning
Following the work (Hu et al., 2024), we evaluate the case-based and rule-based reasoning of Transformer and our FANformer on two tasks, including: (1) Modular addition: $c=(a+b)\mod 113$ with $a,b\in[0,112]$; (2) Linear regression: $c=a+2b+3$ with $a,b\in[0,99]$. We finetune pretrained LLMs, i.e., OLMo-1B and FANformer-1B, on each task dataset for 500 epochs and their performance is measured via the Leave-Square-Out method (10 samples per test point).
@@ Results.
As shown in Figure 10 of Appendix, both Transformer and our FANformer achieve near-perfect accuracy on the training set of modular addition and linear regression tasks, approaching approximately 100%. However, a critical divergence emerges in their performance on the test sets. Specifically, as shown in Figure 5, Transformer exhibits a pronounced failure to generalize, resulting in a "black hole" in the center of the figure, indicating that its accuracy on the test dataset drops to nearly zero. This observation is consistent with the findings reported in the work  (Hu et al., 2024). In contrast, FANformer demonstrates a marked improvement in addressing the "hole" issue. In the linear regression and modular addition tasks, there is no obvious hole observed, further corroborating the hypothesis that, relative to the Transformer-based model, FANformer possesses a stronger tendency to learn underlying rules, thereby achieving superior generalization performance.
4.4.4 LLMs’ Proficiency in Logical Reasoning
Setup.
Following the work (Wang et al., 2024), we adopt the ULogic dataset to systematically evaluate LLMs on their ability to capture underlying inferential logic, where ULogic is constructed after these LLM’s training data cutoff, which prevents data contamination (Dong et al., 2024c). We leverage the Law of Non-Contradiction (Priest et al., 2006), each rule is paired with a negated-conclusion variant; a response is correct only if the model accepts the original and rejects its flip. We evaluate FANformer-1B, OLMo-1B, Qwen2.5-1.5B, and GPT-4 on the two most challenging levels of ULogic for stress-testing.
Table 3: Average performance of different LLMs on ULogic.
Model Acc (%)
GPT-4 65.1
OLMo-1B 0.0
Qwen2.5-1.5B 7.1
FANformer-1B 38.2
Results.
Table 3 summarizes the results on ULogic dataset, with an illustrative case presented in Figure 13 of Appendix. FANformer-1B substantially outperforms both Qwen2.5-1.5B and OLMo-1B: OLMo-1B naively labels every rule as True, and Qwen2.5-1.5B yields largely contradictory answers, whereas FANformer-1B can affirm the original rule and reject its negated counterpart. These results highlight FANformer’s superiority in logic reasoning and underscore the architectural advantages conferred by the periodicity-aware design of FANformer.
4.4.5 Representational Capacity
Figure 6: Representational capacity of MLP, Transformer, and FANformer across various layer depths.
Setup.
We explain the advantage of FANformer architecture from another perspective: the model’s ability to learn complex functions. A larger Lipschitz constant $L$ is frequently linked to greater representational capacity (Latorre et al., 2020; Bartlett et al., 2017), as it enables neural networks to approximate more intricate functional mappings by loosening smoothness constraints in parameter space. This relationship stems from the theoretical framework where less restrictive Lipschitz conditions permit sharper decision boundaries and richer expressive power. Formally, $L$ satisfies:
$\forall x,y\in\mathbb{R}^{n},\quad\lVert f_{\text{model}}(x)-f_{\text{model}}(y)\rVert\leq L\lVert x-y\rVert.$ (9)
Results.
We evaluate the representational capacity of MLP, Transformer, and FANformer across various layer depths via computing their Lipschitz constant $L$. Although representational capacity rises for all three architectures as additional layers are stacked, the growth rates diverge substantially. FANformer exhibits the steepest trajectory and attains the largest Lipschitz constant $L$ at each depth, consistently surpassing Transformer by 27.3% in deeper configurations. These findings indicate that FANformer has markedly richer expressiveness and a superior ability to model complex functions.
@@ 5 Related Work
Large Language Models
@@
The rapid advancement of LLMs has revolutionized NLP and AI research (Radford, 2018; Dubey et al., 2024; DeepSeek-AI et al., 2025). The emergence of GPT-3 (Brown et al., 2020), with 175B parameters, showcased remarkable few-shot prompting abilities, suggesting that scaling laws (Kaplan et al., 2020) could unlock emergent capabilities. Recent notable LLMs like PaLM (Chowdhery et al., 2023), LLaMA (Touvron et al., 2023b), GPT-4 (Chowdhery et al., 2023), and DeepSeek (Bi et al., 2024) further pushed the boundaries of model size and performance. Moreover, the open-source release of OLMo (Groeneveld et al., 2024) has provided valuable resources for the community, enabling more accessible training of LLMs.
Advances in Transformer Architecture
Recent advancements in Transformer architecture focus on overcoming two main limitations: computational inefficiency in long-context processing and limited expressiveness of attention mechanisms. To address long-context inefficiencies, techniques such as sparsity (Child et al., 2019) and local attention (Beltagy et al., 2020), along with innovations in query mechanisms (Shazeer, 2019; Ainslie et al., 2023) and inference acceleration (DeepSeek-AI et al., 2024b), have been proposed. Hardware-level optimizations like FlashAttention (Dao et al., 2022) further reduce GPU memory access overhead. To enhance the expressiveness of attention mechanisms, methods like Probabilistic Attention Keys (Nguyen et al., 2022) and Selective Attention (Leviathan et al., 2024) improve semantic relationship capture and refine attention by suppressing irrelevant features. Additionally, approaches like Differential Transformer address attention noise in long contexts (Ye et al., 2024). Different from previous work, we improve language modeling by addressing the challenge of periodicity modeling in Transformers, which can seamlessly incorporate the aforementioned works for revising the attention mechanism, as demonstrated in the derivation provided in Appendix I.
Fourier-based Networks
@@
Previous research on Fourier-based Networks was aimed at solving some domain-specific applications (Zuo and Cai, 2005; Tan, 2006; Chen et al., 2022; Li et al., 2021). Some studies specifically explored the use of sinusoidal activations (e.g., cosine (Silvescu, 1999) (Ngom and Marin, 2021) or sine (Parascandolo et al., 2016; Sitzmann et al., 2020)) to approximate periodic patterns (Liu, 2013). However, these approaches lacked generalizability beyond narrow domains due to rigid frequency parameterization and limited scalability (Uteuliyeva et al., 2020; Liu et al., 2020). Recent work (Dong et al., 2024a) addresses these problems using FAN to introduce Fourier Principle into neural networks, but its adaptation to LLMs remains an open challenge. FNet (Lee-Thorp et al., 2022) replaces self-attention with Fourier Transform to achieve linear complexity, but it sacrifices the performance of LMs. In contrast, we employ effective periodicity modeling to improve LLMs.
@@ 6 Conclusion
We propose FANformer, a novel LLM architecture that enhances learning efficiency by adapting Fourier Analysis Network into attention mechanism for effective periodicity modeling. Experiments show that FANformer outperforms Transformer when scaling model parameters and training tokens, achieving better performance with 31% fewer parameters and 20% fewer tokens. Pretrained FANformer-1B surpasses open-source LLMs of comparable size or training scale on various downstream tasks. The discovery of FANformer’s enhanced scalability, learning efficiency, rule-based learning advantages, and representational capacity suggests potential pathways for developing more efficient and high-performance language models.
@@ Appendix A Training Loss Curves of OLMO and FANformer
Appendix B Preliminary Knowledge
Fourier Analysis Network (FAN) [Dong et al., 2024a] enhances neural networks by introducing Fourier principles for effective periodicity modeling. The core component of FAN is its layer design, which combines periodic basis functions with standard linear transformations. Given a input $\mathbf{X}$, the FAN layer is defined as:
$\text{FANLayer}(\mathbf{X})=[\cos(W_{p}\mathbf{X})\|\sin(W_{p}\mathbf{X})\|\sigma(W_{\bar{p}}\mathbf{X}+B_{\bar{p}})]$ (10)
[…]
Appendix C Effect of hyperparameter $p$
We systematically investigate the impact of hyperparameter $p$, which controls the proportion of periodicity modeling in FANformer, on model performance across its value range. The experimental results from the 1B-scale FANformer (as shown in Figure 8) demonstrate that our model exhibits strong robustness in terms of training loss and downstream task accuracy, with relatively small performance fluctuations. Furthermore, regardless of the variation in p values, FANformer consistently outperforms the standard Transformer (horizontal baseline). Analysis of experimental results from models of different scales (300M, 1B, 3B) (as shown in Figure 9) reveals a clear trend: larger models tend to exhibit higher optimal p values. This observation suggests that more powerful FANformers are better equipped to extract more intricate latent periodicity features.
@@ Appendix C Effect of hyperparameter $p$
Appendix D Instruction Following with SFT
@@ D.2 Evaluation Setup
Appendix E Computational Cost Analysis
We analyze the computational overhead of FANformer compared to standard Transformer under two settings: (1) when the number of parameters is matched, the computational cost remains unchanged; (2) when the FFN dimensions remain fixed, the increase in cost is negligible.
To ensure a fair comparison, we maintain the same number of parameters in FANformer as in the original Transformer by varying the FFN’s intermediate dimension $d_{f}$ during our experiments. Under this configuration, the computational cost remains equivalent between FANformer (same parameter) and Transformer. We further investigate the setting where FFN dimensions remain unchanged (Table 5). Even for FANformer (same dimension), the computational cost increases only slightly compared to the Transformer. Specifically, the ratio of additional computation can be expressed as:
$\frac{1.5LSD^{2}}{L\times(24SD^{2}+4S^{2}D)+2SDV}=\frac{1.5D}{24D+4S}\leq\frac{1.5D}{24D}=0.0625,$ (11)
where $L$ is the number of layers, $S$ is the sequence length, $D$ is the hidden dimension, and $V$ is the vocabulary size. In practice, the actual additional overhead is much lower than $0.0625$, as the term scales inversely with sequence length. This means the overhead diminishes rapidly as sequences grow longer. We summarize the detailed comparison of computational costs between Transformer and FANformer in Table 5.
Table 5: Computational cost of Transformer vs. FANformer ($p=0.25$).
Model Self-Attention/ATF FLOPs FFN FLOPs Output FLOPs Total FLOPs
Transformer $L\times(8SD^{2}+4S^{2}D)$ $L\times 16SD^{2}$ $2SDV$ $L\times(24SD^{2}+4S^{2}D)+2SDV$
FANformer (Same Param) $L\times(8SD^{2}+4S^{2}D+1.5SD^{2})$ $L\times(16SD^{2}-1.5SD^{2})$ $2SDV$ $L\times(24SD^{2}+4S^{2}D)+2SDV$
FANformer (Same Dim) $L\times(8SD^{2}+4S^{2}D+1.5SD^{2})$ $L\times 16SD^{2}$ $2SDV$ $L\times(25.5SD^{2}+4S^{2}D)+2SDV$
Appendix F Detailed Results of Ablation Study for Section 4.4.1
In ablation study, we report the average results across various tasks on V2 Validation Sets, V3 Validation Sets, and Downstream tasks, with the specific tasks detailed in Section J.3. The complete results are detailed in Table 6 and Table 7.